OpenStudy (anonymous):

integrate: (x^2+4)/(x^2-4)

OpenStudy (turingtest):

first use long division, then factor the remaining numerator and use partial fractions

OpenStudy (anonymous):

after long division i got 1 remainder 8

OpenStudy (turingtest):

ok, so that would be 8/(x^2-4) now you can factor the denominator (*oh I said numerator last time, my bad..) and use partial fractions

OpenStudy (anonymous):

factoring gives you (x+2)(x-2)

OpenStudy (turingtest):

can you perform partial fractions on that?

OpenStudy (anonymous):

no

OpenStudy (turingtest):

not familiar with the technique?

OpenStudy (anonymous):

there's only one number in the numerator

OpenStudy (anonymous):

i've done it before but don't remember much

OpenStudy (turingtest):

\[{8\over(x-2)(x+2)}={A\over x-2}+{B\over x+2}\]looking familiar yet?

OpenStudy (turingtest):

\[{8\over(x-2)(x+2)}={A\over x-2}+{B\over x+2}\implies A(x+2)+B(x-2)=8\]now we choose some values for x that will let us know what A and B are

OpenStudy (anonymous):

ok let me try to work this. i remember now

OpenStudy (anonymous):

thank you

OpenStudy (turingtest):

welcome :)

OpenStudy (anonymous):

2 nd -2 will make it zero

OpenStudy (turingtest):

so A=? B=?

OpenStudy (anonymous):

B=-2 and A=2

OpenStudy (turingtest):

yep :) so our integral is now...?

OpenStudy (anonymous):

(-2/x+2) + (2/x-2) +1

OpenStudy (turingtest):

yes, and you can integrate that, I'm pretty sure good job !

OpenStudy (anonymous):

2ln(x-2/x+2) +x

OpenStudy (anonymous):

i forgot my constant and of course +C